The owner of D. Donuts Ltd. wants to buy a new machine for his bakery. He is looking at two possible machines, Machine A and Machine B. On the one hand, Machine A will initially cost $ 100, its operating cost will be $ 10 per year and is expected to last 2 years. On the other hand, Machine B will cost $ 140 with an operating cost of $ 8 per year, and will run out after 3 years. Both machines will be amortized at an ACC rate of 25%. If the ERR is 10% and the tax rate is 42%, which machine should the owner buy?
1
Expert's answer
2020-03-18T10:02:28-0400


1) By car A

Calculate cash flow by year

"Cf1=100+100\times0,42+0,25\times100+10=177"

"Cf2=10+100\times0,25=35"

"Cf3=0,25\times100=25"

"Cf4=0,25\times100=25"

Total cash flow: "262=177+35+25+25"

Calculate the discounted cash flow:

"DCf1=\frac{177} {(1+0.1)^1}=161"

"DCf2=\frac{35}{(1+0.1)^2}=29"

"DCf3=\frac{25}{(1+0.1)^3}=19"

"DCf4=\frac{25}{(1+0.1)^4}=17"

Total discounted cash flow: "226=161+29+19+17"

"NPVa=226-100=126"

2)By car B

Calculate cash flow by year

"Cf1=140+140\times0,42+0,25\times140+8=242"

"Cf2=8+140\times0,25=43"

"Cf3=8+140\times0,25=43"

"Cf4=140\times0,25=35"

Total cash flow: "363=242+43+43+35"

Calculate the discounted cash flow:

"DCf1=\frac{242}{(1+0.1)^1}=220"

"DCf2=\frac{43}{(1+0.1)^2}=36"

"DCf3=\frac{43}{(1+0.1)^3}=32"

"DCf4=\frac{35}{(1+0.1)^4}=24"

Total discounted cash flow: "220+36+32+24=312"

"NPVb=312-140=172"

NPVb more NPVa, owner buys car b